OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, -3) with a slope of -3? a.3x + y = 12 b.3x + y = -12 c.3x - y = 12 d.3x - y = -12

OpenStudy (tkhunny):

Please show your work. Point-Slope Form is this: (y-k) = m(x-h) We have a point: (5,-3), thus (y+3) = m(x-5) We have the slope: -3, thus (y+3) = -3(x-5) Now, for standard form: y + 3 = -3(x-5) Expand parentheses y + 3 = -3x + 15 Add 3x to get variables on the left. 3x + y + 3 = 15 Subtract 3 to get constants on the right. 3x + y = 12 -- I think we're done! Just one step at a time. No need to make the whole trip at the beginning!